Rainbows, stunning as they can be, are a result of diffraction of light through raindrops. Calculate the energy of a photon of red light in the rainbow with a wavelength of 665.0 nm and the energy of a photon of purple light in the rainbow with a wavelength of 425.0 nm

(A). Determine the energy of a photon of each wavelength.

Photon of Red Light λ = 665.0 nm

   Photon of Blue Light λ = 425.0 nm
Answer #1

1) for 655.0 nm

Given:
lambda = 665.0 nm = 6.650*10^-7 m

use:
E = h*c/lambda
=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(6.650*10^-7 m)
= 2.989*10^-19 J
Answer: 2.989*10^-19 J

2) for 425.0 nm

Given:
lambda = 425.0 nm = 4.250*10^-7 m

use:
E = h*c/lambda
=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(4.250*10^-7 m)
= 4.677*10^-19 J
Answer: 4.677*10^-19 J